# Chain-of-Retrieval Augmented Generation

### v2

@@ Abstract
(a) Test-time scaling behavior of CoRAG.
(b) An example of CoRAG in action.
@@ 1 Introduction
To break the bottleneck of retrieval quality, we propose a framework that dynamically retrieves relevant information and plans subsequent retrieval steps based on the current state. By adjusting the number of retrieval steps at test time, our model can explore various aspects of the query and experiment with different query rewriting strategies when the retriever does not yield useful information. This paradigm mirrors the human problem solving process, where we iteratively seek information to address complex questions.
@@
On the KILT benchmark [25], which encompasses a more diverse array of tasks, new state-of-the-art scores are achieves on the hidden test set for nearly all tasks. Additionally, we uncover that CoRAG exhibits varied scaling behaviors across different task types. For datasets such as NQ [17], where state-of-the-art retrievers already achieve high recall, the benefits of test-time scaling are often marginal. This suggests the potential for dynamically allocating test-time compute based on the complexity of the query and the quality of the retriever. Upon further analysis, we find that CoRAG can effectively decompose complex queries and perform flexible query reformulation to improve the quality of the generated responses. It also shows robustness against retrievers of varying quality. We posit that CoRAG represents a promising avenue for future research in the RAG domain, with the potential to mitigate hallucination in model-generated content. Our code is available at https://github.com/microsoft/LMOps/tree/main/corag.
@@ 2 Related Work
To mitigate this constraint, contemporary research has extended the conventional paradigm of a single retrieval step followed by generation, advancing to multi-step iterative retrieval and generation [5]. FLARE [12] prompts an LLM to actively determine when and what to retrieve during the generation process. ITER-RETGEN [28] proposes to interleave retrieval-augmented generation with generation-augmented retrieval, demonstrating enhancements in multi-hop QA tasks. Similarly, IRCoT [31] employs a chain-of-thought methodology, which recursively refines the reasoning thought for subsequent retrieval steps. Self-RAG [1] empowers LLMs to adaptively retrieve, generate, and critique through self-reflection, thus improving factual accuracy and citation precision in open-domain QA and long-form generation tasks. Auto-RAG [38] utilizes heuristic rules and exact answer matching to construct intermediate retrieval steps, yet its performance remains significantly below that of state-of-the-art models. In this study, rather than exclusively on few-shot prompting or distillation from proprietary models, we propose a novel approach to explicitly train LLMs to iteratively retrieve and reason over relevant information.
@@
In the realm of RAG, test-time compute can be increased by retrieving more documents or performing additional retrieval steps. LongRAG [13] posits that RAG performance can be enhanced by integrating long-context LLMs with more retrieved documents. In contrast, IterDRAG [39] empirically examines the test-time scaling law through few-shot prompting and iterative retrieval for up to $5$M tokens. A concurrent work Search-o1 [20] combines the open-source QwQ model [35] with active search from Bing, achieving competitive results on knowledge-intensive tasks. Our work extends the study of test-time scaling in RAG to a targeted fine-tuning paradigm under diverse decoding strategies.
@@ 3 Methodology
The CoRAG framework is illustrated in Figure 1. In this section, we describe the key components of CoRAG, including retrieval chain generation through rejection sampling, model training with augmented datasets, and strategies for scaling test-time compute.
@@ 3.2 Training
Each training instance in the augmented dataset is represented as a tuple $(Q,A,Q_{1:L},A_{1:L})$, accompanied by the corresponding top-$k$ retrieved documents for the query $Q$ and each sub-query. We fine-tune an LLM on the augmented dataset using the standard next-token prediction objective within a multi-task learning framework.
@@ 3.3 Test-time Scaling
Best-of-$N$ Sampling This method involves sampling $N$ retrieval chains with a temperature $0.7$, subsequently selecting the best chain to generate the final answer. As the ground truth answer is not available at test time, we instead calculate the conditional log-likelihood of “No relevant information found” as a penalty score for each chain. The retrieval chain with the lowest penalty score is chosen.
@@ 4.1 Setup
Model Training We conduct full-parameter fine-tuning on the augmented datasets, initializing from the Llama-3.1-8B-Instruct checkpoint. Two separate models are trained: one for the multi-hop QA datasets and another for the KILT benchmark. The compiled multi-hop QA dataset comprises $125$k training instances, whereas the KILT benchmark includes $660$k instances after sub-sampling. The model is fine-tuned for $1$ epoch with a maximum sequence length of $3$k tokens. For the KILT benchmark, we fine-tune an E5-Mistral retriever [33] and a RankLLaMA re-ranker [22] on the respective training set to boost the ranking quality.
@@ 4.2 Main Results
$3$-shot Llama-3.1-8B-Inst. 27.6 32.1 20.8 28.8 17.6 21.3 3.4 9.7
$3$-shot GPT-4o 39.5 47.3 38.2 51.2 49.6 61.5 15.8 27.2
@@
$3$-shot Llama-3.1-8B-Inst. 30.7 39.9 34.1 46.6 28.0 37.3 7.7 15.4
$3$-shot GPT-4o 49.0 56.2 45.8 59.4 53.6 63.8 15.7 25.8
@@
    $\triangleright$ $L$=$1$, greedy 56.5 62.3 50.1 63.2 37.6 51.4 18.6 29.3
    $\triangleright$ $L$=$6$, greedy 70.6 75.5 54.4 67.5 48.0 63.5 27.7 38.5
    $\triangleright$ $L$=$6$, best-of-$4$ 71.7 76.5 55.3 68.5 51.2 63.1 28.1 39.7
    $\triangleright$ $L$=$6$, tree search 71.7 76.4 55.8 69.0 48.8 64.4 29.0 40.3
    $\triangleright$ $L$=$10$, best-of-$8$ 72.5 77.3 56.3 69.8 54.4 68.3 30.9 42.4
@@ 4.3 Scaling Test-Time Compute
In alignment with OpenAI o1 [11], our model allows for scaling test-time compute to potentially achieve better performance without updating model weights. There are multiple ways to control the test-time compute. In Figure 2, we concentrate on two factors: the retrieval chain length $L$ and the number of sampled chains $N$ for best-of-$N$ sampling. Greedy decoding is a special instance of best-of-$N$ sampling with $N=1$ and the temperature set to 00.
@@
The Pareto frontier between the EM score and token consumption approximately follows a log-linear trajectory for up to $128$k tokens, although the scaling behavior varies across different datasets. This observation assists practitioners in making informed decisions regarding the allocation of test-time compute based on the quality requirements. It is important to note that we make several simplifications in this scaling study, such as treating the prompt tokens equivalently to the generated tokens and ignoring the retrieval costs. A more rigorous analysis could take these factors into account.
@@ 5 Analysis
Table 3: Ablation study results. “Iterative training” employs a trained CoRAG model for another round of rejection sampling. “Weak-to-strong Generalization” utilizes weaker LLMs for retrieval chain generation while using stronger LLMs (Llama-3.1-8B-Inst.) for training. “Different Retrievers” replaces the text retriever at test time.
@@ 5.2 Robustness and Generalization
Different Retrievers We further investigate the influence of various text retrievers at test time. Instead of using the E5-large dense retriever, we substitute it with two weaker alternatives in a plug-and-play fashion: E5-base and BM25. Across all datasets, we observe consistent performance gains when investing more test-time compute, although stronger retrievers continue to outperform in terms of absolute performance. Orthogonal efforts to enhance the quality of text retrievers are likely to further boost the performance of CoRAG.
@@
The results in Table 3 demonstrate that utilizing Llama-3B achieves very close performance compared to the 8B model, whereas Llama-1B exhibits a noticeable performance drop. Manual inspection reveals that the $1$B model frequently struggles to follow the given instructions, resulting in sub-optimal retrieval chains. Employing weaker LLMs also lowers the barrier to adopting more computationally expensive tree search strategies during data generation, which show great potential in mathematical reasoning tasks [6].
@@ 5.3 Does Chain-of-Retrieval Always Help?
Multi-hop QA datasets are specifically designed to evaluate complex reasoning capabilities and are expected to benefit from the chain-of-retrieval mechanism. Table  1 presents empirical evidence supporting this assertion. In contrast, for tasks that a single retrieval step is typically sufficient, the advantage tends to be marginal, as demonstrated in Figure  3. Datasets such as NQ [17] and TriviaQA [15] are known for their (mostly) single-hop nature. This phenomenon implies that decoding strategies should be adaptive based on the complexity of the query. Additional results on the full KILT benchmark are listed in Appendix Table 5, where similar observations for other task types also hold.
@@ 5.4 Learning to Stop at Test Time
During the training phase, an additional loss term is added for the stop prediction task. The target output is “Yes” if the current retrieval chain encompasses the prefix that maximizes the likelihood of the final answer, and “No” otherwise. The associated prompt template is in Appendix Section C.
In Figure  4, we illustrate how the performance varies along with the token consumption on the MuSiQue dataset. While early stopping can save some amount of token quota, it comes at the cost of performance degradation. The optimal configuration depends on the dataset characteristics and the quality expectations.
@@ Appendix A Implementation Details
Rejection Sampling For each training instance, we sample up to $16$ retrieval chains, with the maximum length randomly selected from the interval $[1,5]$. The sampling temperature is set to $0.7$ for sub-query generation and 00 for sub-answer generation. Chain generation is terminated if the sub-answer matches the correct answer or if the average conditional log-likelihood of the correct answer exceeds $-0.05$. For each sub-query, we utilize the E5-large retriever  [^1] https://huggingface.co/intfloat/e5-large-v2 to retrieve the top-$5$ most relevant documents from the KILT version of the Wikipedia corpus [23]. This corpus comprises $36$ million passages.
@@
Multi-Hop QA Training Hyperparameters The training set is the union of the 2WikiMultihopQA, HotpotQA, and MuSiQue datasets, comprising a total of $125$k samples. The Bamboogle dataset, consisting of only $125$ questions, is reserved for evaluation only. Additional hyperparameters are detailed in Table 4. To balance the three loss terms in Section 3.2, we set a sample ratio of $0.2$ for both the sub-query and sub-answer generation tasks; this ratio is also applied to the KILT training.
KILT Training Hyperparameters We utilize the official training set of the KILT benchmark, omitting the ELI5 and WoW datasets due to the lack of reliable evaluation metrics. To balance the task distribution, we only select $100$k samples for large datasets like T-REx and Zero-Shot RE. In accordance with the benchmark’s guidelines, we also add $100$k samples from the BLINK dataset for entity linking.
@@
Decoding Strategies In the context of best-of-$N$ sampling, the temperature is set to $0.7$ for sub-query generation. For sub-answer generation and final answer prediction, the temperature is always set to 00 across all decoding strategies. Regarding tree search, we set the expansion size to $4$ and the number of rollouts to $2$. Given that tree search incurs a significantly higher token consumption compared to other decoding strategies, we limit the rollouts to a maximum of $2$ steps for each expansion. To avoid the model from generating repetitive sub-queries endlessly, any generated sub-query identical to previous ones is discarded.
@@ Appendix B Additional Results
    $\triangleright$ $L$=$1$, greedy 90.4 86.0 76.8 87.0 82.1 62.5 56.4 88.4 91.4
    $\triangleright$ $L$=$6$, greedy 92.7 87.4 75.8 86.6 83.8 63.2 59.1 88.6 93.8
    $\triangleright$ $L$=$6$, best-of-$4$ 92.5 87.4 75.8 86.3 83.5 62.6 59.6 88.7 93.9
    $\triangleright$ $L$=$6$, tree search 91.8 86.8 75.5 86.4 83.0 62.4 59.9 88.9 93.9
@@
Figure 5: Scaling rejection sampling compute for training data generation. We vary the number of sampled chains from $4$ to $16$ while maintaining all other hyperparameters fixed.
@@
Appendix C Prompts

### v3

@@ Abstract
Figure 1: (a) Test-time scaling behavior of CoRAG. Increased token budget leads to consistent performance improvements. (b) An example of CoRAG on the MuSiQue dataset. It learns to decompose the complex query and conduct query reformulation when encountering a retrieval failure.
@@ 1 Introduction
To break the bottleneck of retrieval quality, we propose a framework that dynamically retrieves relevant information and plans subsequent retrieval steps based on the current state. By adjusting the number of retrieval steps at test time, our model can explore various aspects of the query and experiment with different query rewriting strategies when the retriever does not yield useful information. This paradigm mirrors the human problem solving process, where we iteratively seek information to address complex questions. An example is illustrated in Figure 1.
@@
On the KILT benchmark [27], which encompasses a more diverse array of tasks, new state-of-the-art scores are achieves on the hidden test set for nearly all tasks. Additionally, we uncover that CoRAG exhibits varied scaling behaviors across different task types. For datasets such as NQ [19], where state-of-the-art retrievers already achieve high recall, the benefits of test-time scaling are often marginal. This suggests the potential for dynamically allocating test-time compute based on the complexity of the query and the quality of the retriever. Upon further analysis, we find that CoRAG can effectively decompose complex queries and perform flexible query reformulation to improve the quality of the generated responses. It also shows robustness against retrievers of varying quality. We posit that CoRAG represents a promising avenue for future research in the RAG domain, with the potential to mitigate hallucination in model-generated content. Our code, data and trained models are available at https://github.com/microsoft/LMOps/tree/main/corag.
@@ 2 Related Work
To mitigate this constraint, contemporary research has extended the conventional paradigm of a single retrieval step followed by generation, advancing to multi-step iterative retrieval and generation [6]. FLARE [13] prompts an LLM to actively determine when and what to retrieve during the generation process. ITER-RETGEN [30] proposes to interleave retrieval-augmented generation with generation-augmented retrieval, demonstrating enhancements in multi-hop QA tasks. Similarly, IRCoT [33] employs a chain-of-thought methodology, which recursively refines the reasoning thought for subsequent retrieval steps. Self-RAG [1] empowers LLMs to adaptively retrieve, generate, and critique through self-reflection, thus improving factual accuracy and citation precision in open-domain QA and long-form generation tasks. Auto-RAG [41] utilizes heuristic rules and exact answer matching to construct intermediate retrieval steps, yet its performance remains significantly below that of state-of-the-art models. AQA [3] learns to reformulate questions using reinforcement learning but only focuses on single-hop QA tasks. In this study, rather than exclusively on few-shot prompting or distillation from proprietary models, we propose a novel approach to explicitly train LLMs to iteratively retrieve and reason over relevant information.
@@
In the realm of RAG, test-time compute can be increased by retrieving more documents or performing additional retrieval steps. LongRAG [14] posits that RAG performance can be enhanced by integrating long-context LLMs with more retrieved documents. In contrast, IterDRAG [42] empirically examines the test-time scaling law through few-shot prompting and iterative retrieval for up to 55M tokens. Search-o1 [22] combines the open-source QwQ model [37] with active search from Bing, achieving competitive results on knowledge-intensive tasks. Concurrent works such as Search-R1 [15] train LLMs to use retrieval as a tool via reinforcement learning. Our work extends the study of test-time scaling in RAG to a targeted fine-tuning paradigm under diverse decoding strategies.
@@ 3 Methodology
The CoRAG framework is illustrated in Figure 2. The “Current State” denotes the input context and instructions provided to the LLM, while the “Next Action” refers to the LLM output responding to the given instruction. In this section, we describe the key components of CoRAG, including retrieval chain generation through rejection sampling, model training with augmented datasets, and strategies for scaling test-time compute.
@@ 3.2 Training
Each training instance in the augmented dataset is represented as a tuple $(Q,A,Q_{1:L},A_{1:L})$, accompanied by the corresponding top-$k$ retrieved documents for the query $Q$ and each sub-query. We fine-tune an LLM on the augmented dataset using the standard next-token prediction objective within a unified multi-task learning framework.
@@ 3.3 Test-time Scaling
Best-of-NN Sampling This method involves sampling $N$ retrieval chains with a temperature $0.7$, subsequently selecting the best chain to generate the final answer. As the ground truth answer is not available at test time, we instead calculate the conditional log-likelihood of “No relevant information found” as a penalty score for each chain. The retrieval chain with the lowest penalty score is chosen.
@@ 4.1 Setup
Note that while HotpotQA and MuSiQue maintain public leaderboards, these adopt either a simplified reading comprehension setting or an abstract-only retrieval configuration. Consequently, the leaderboard results are not directly comparable to our open-domain QA evaluation setting.
Model Training We conduct full-parameter fine-tuning on the augmented datasets, initializing from the Llama-3.1-8B-Instruct checkpoint. Two separate models are trained: one for the multi-hop QA datasets and another for the KILT benchmark. The compiled multi-hop QA dataset comprises 125125k training instances, whereas the KILT benchmark includes 660660k instances after sub-sampling. The model is fine-tuned for $1$ epoch with a maximum sequence length of 33k tokens. For the KILT benchmark, we fine-tune an E5-Mistral retriever [35] and a RankLLaMA re-ranker [24] on the respective training set to boost the ranking quality.
@@ 4.2 Main Results
33-shot Llama-3.1-8B-Inst. 27.6 32.1 20.8 28.8 17.6 21.3 3.4 9.7
33-shot GPT-4o 39.5 47.3 38.2 51.2 49.6 61.5 15.8 27.2
@@
33-shot Llama-3.1-8B-Inst. 30.7 39.9 34.1 46.6 28.0 37.3 7.7 15.4
33-shot GPT-4o 49.0 56.2 45.8 59.4 53.6 63.8 15.7 25.8
@@
   $\triangleright$ LL=11, greedy 56.5 62.3 50.1 63.2 37.6 51.4 18.6 29.3
   $\triangleright$ LL=66, greedy 70.6 75.5 54.4 67.5 48.0 63.5 27.7 38.5
   $\triangleright$ LL=66, best-of-44 71.7 76.5 55.3 68.5 51.2 63.1 28.1 39.7
   $\triangleright$ LL=66, tree search 71.7 76.4 55.8 69.0 48.8 64.4 29.0 40.3
   $\triangleright$ LL=$10$, best-of-88 72.5 77.3 56.3 69.8 54.4 68.3 30.9 42.4
@@ 4.3 Scaling Test-Time Compute
In alignment with OpenAI o1 [12], our model allows for scaling test-time compute to potentially achieve better performance without updating model weights. There are multiple ways to control the test-time compute. In Figure 3, we concentrate on two factors: the retrieval chain length $L$ and the number of sampled chains $N$ for best-of-$N$ sampling. Greedy decoding is a special instance of best-of-$N$ sampling with $N=1$ and the temperature set to 0.
@@
The Pareto frontier between the EM score and token consumption approximately follows a log-linear trajectory for up to 128128k tokens, although the scaling behavior varies across different datasets. This observation assists practitioners in making informed decisions regarding the allocation of test-time compute based on the quality requirements. It is important to note that we make several simplifications in this scaling study, such as treating the prompt tokens equivalently to the generated tokens and ignoring the retrieval costs. A more rigorous analysis could take these factors into account.
@@ 5 Analysis
Table 3: Ablation study results. “Iterative training” employs a trained CoRAG model for another round of rejection sampling. “Distill from GPT-4o” leverages the GPT-4o model to generate retrieval chains. “Weak-to-strong Generalization” utilizes weaker LLMs for retrieval chain generation while using stronger LLMs (Llama-3.1-8B-Inst.) for training. “Different Retrievers” replaces the text retriever at test time.
@@ 5.2 Robustness and Generalization
Different Retrievers We further investigate the influence of various text retrievers at test time. Instead of using the E5-large dense retriever, we substitute it with two weaker alternatives in a plug-and-play fashion: E5-base and BM25. Across all datasets, we observe consistent performance gains when investing more test-time compute, although stronger retrievers continue to outperform in terms of absolute performance. Improvements to text retriever quality represent an orthogonal dimension that can further amplify CoRAG’s performance gains.
@@
The results in Table 3 demonstrate that utilizing Llama-3B achieves very close performance compared to the 8B model, whereas Llama-1B exhibits a noticeable performance drop. Manual inspection reveals that the 11B model frequently struggles to follow the given instructions, resulting in sub-optimal retrieval chains. Employing weaker LLMs also lowers the barrier to adopting more computationally expensive tree search strategies during data generation, which show great potential in mathematical reasoning tasks [7]. In contrast, distilling from a stronger model like GPT-4o yields a further performance boost, indicating that the quality of the retrieval chains is crucial for the final performance.
@@ 5.3 Does Chain-of-Retrieval Always Help?
Multi-hop QA datasets are specifically designed to evaluate complex reasoning capabilities and are expected to benefit from the chain-of-retrieval mechanism. Table  1 presents empirical evidence supporting this assertion. In contrast, for tasks that a single retrieval step is typically sufficient, the advantage tends to be marginal, as demonstrated in Figure  4. Datasets such as NQ [19] and TriviaQA [17] are known for their (mostly) single-hop nature. This phenomenon implies that decoding strategies should be adaptive based on the complexity of the query. Additional results on the full KILT benchmark are listed in Appendix Table 7, where similar observations for other task types also hold.
@@ 5.4 Learning to Stop at Test Time
During the training phase, an additional loss term is added for the stop prediction task. The target output is “Yes” if the current retrieval chain encompasses the prefix that maximizes the likelihood of the final answer, and “No” otherwise. The associated prompt template is in Appendix Section D.
In Figure  5, we illustrate how the performance varies along with the token consumption on the MuSiQue dataset. While early stopping can save some amount of token quota, it comes at the cost of performance degradation. The optimal configuration depends on the dataset characteristics and the quality expectations.
5.5 Does CoRAG Learn to Retrieve Better?
[…]
Table 4: Retrieval recall comparison between standard retrieval and CoRAG across multi-hop QA datasets.
R@10 R@20 R@100
[…]
   w/ CoRAG 72.1 76.7 84.3
[…]
   w/ CoRAG 81.4 84.8 88.8
[…]
   w/ CoRAG 59.2 68.0 75.2
[…]
   w/ CoRAG 47.1 54.6 68.4
The results in Table  4 demonstrate that CoRAG consistently improves recall across all datasets and recall thresholds. The improvements are particularly pronounced on more challenging datasets like MuSiQue and Bamboogle, where single-step retrieval struggles most. This indicates that CoRAG’s iterative query reformulation and decomposition strategy effectively addresses the limitations of traditional dense retrieval, enabling the model to gather more relevant information through multiple retrieval steps.
@@ 6 Conclusion
7 Limitations and Broader Impacts
This study primarily investigates RAG tasks characterized by short and easy-to-verify answers, such as multi-hop QA and entity linking. However, real-world applications often necessitate addressing more complex tasks that demand generating long-form outputs. A significant challenge in long-form generation lies in the absence of robust evaluation metrics within the current research landscape.
Regarding broader impacts, the proposed framework aims to improve the factuality and groundedness of language model outputs. It is anticipated that this work can facilitate more efficient and effective information retrieval for users. Nevertheless, the inherent risk of hallucination persists and warrants careful monitoring in practical deployments.
@@ Appendix A Implementation Details
Rejection Sampling For each training instance, we sample up to $16$ retrieval chains, with the maximum length randomly selected from the interval $[1,5]$. The sampling temperature is set to $0.7$ for sub-query generation and 0 for sub-answer generation. Chain generation is terminated if the sub-answer matches the correct answer or if the average conditional log-likelihood of the correct answer exceeds 0.05-0.05. For each sub-query, we utilize the E5-large retriever  [^1] https://huggingface.co/intfloat/e5-large-v2 to retrieve the top-$5$ most relevant documents from the KILT version of the Wikipedia corpus [25]. This corpus comprises $36$ million passages.
@@
Table 6: Statistics of the datasets used for multi-hop QA training.
2WikiMultihopQA HotpotQA Bamboogle MuSiQue
# Training Samples $15,000$ $90,447$ - $19,938$
# Validation Samples $12,576$ $7,405$ $125$ $2,417$
Multi-Hop QA Training Hyperparameters The training set is the union of the 2WikiMultihopQA, HotpotQA, and MuSiQue datasets, comprising a total of 125125k samples as shown in Table 6. The Bamboogle dataset, consisting of only $125$ questions, is reserved for evaluation only. Additional hyperparameters are detailed in Table 5. To balance the three loss terms in Section 3.2, we set a sample ratio of $0.2$ for both the sub-query and sub-answer generation tasks; this ratio is also applied to the KILT training.
KILT Training Hyperparameters We utilize the official training set of the KILT benchmark, omitting the ELI5 and WoW datasets due to the lack of reliable evaluation metrics. To balance the task distribution, we only select 100100k samples for large datasets like T-REx and Zero-Shot RE. In accordance with the benchmark’s guidelines, we also add 100100k samples from the BLINK dataset for entity linking.
@@
Decoding Strategies In the context of best-of-$N$ sampling, the temperature is set to $0.7$ for sub-query generation. For sub-answer generation and final answer prediction, the temperature is always set to 0 across all decoding strategies. Regarding tree search, we set the expansion size to $4$ and the number of rollouts to $2$. Given that tree search incurs a significantly higher token consumption compared to other decoding strategies, we limit the rollouts to a maximum of $2$ steps for each expansion. To avoid the model from generating repetitive sub-queries endlessly, any generated sub-query identical to previous ones is discarded.
@@ Appendix B Additional Results
   $\triangleright$ LL=11, greedy 90.4 86.0 76.8 87.0 82.1 62.5 56.4 88.4 91.4
   $\triangleright$ LL=66, greedy 92.7 87.4 75.8 86.6 83.8 63.2 59.1 88.6 93.8
   $\triangleright$ LL=66, best-of-44 92.5 87.4 75.8 86.3 83.5 62.6 59.6 88.7 93.9
   $\triangleright$ LL=66, tree search 91.8 86.8 75.5 86.4 83.0 62.4 59.9 88.9 93.9
@@
Figure 6: Scaling rejection sampling compute for training data generation. We vary the number of sampled chains from 44 to $16$ while maintaining all other hyperparameters fixed.
@@
Figure 7: Scaling test-time compute on multi-hop QA datasets with Llama-3.1-8B-Instruct. No fine-tuning is performed on the model weights.
Scaling Test-Time Compute without Model Fine-Tuning In Figure  7, we present the scaling results on multi-hop QA datasets using the Llama-3.1-8B-Instruct model directly without any fine-tuning. The scaling curves are similar to those observed in Figure 3, but the absolute performance is significantly lower, indicating that targeted fine-tuning is essential for improving the scaling upper bound.
@@
Table 9: Extension to the Qwen3 model families.
2WikiQA HotpotQA Bamboogle MuSiQue
EM F1 EM F1 EM F1 EM F1
[…]
CoRAG-Qwen3-4B (L=6, greedy) 69.3 74.1 51.6 64.2 49.6 62.5 24.0 34.5
[…]
CoRAG-Qwen3-8B (L=6, greedy) 70.0 74.8 52.8 66.0 49.6 63.7 25.2 35.9
Extension to Other Model Families To demonstrate that our CoRAG framework is agnostic to model families and not limited to Llama-based architectures, we conduct experiments using Qwen3-4B and Qwen3-8B [38] models following the same training procedure. As shown in Table 9, CoRAG consistently outperforms the baseline fine-tuned models across all datasets and model sizes, with improvements of over $10$ EM points on average. This validates that the chain-of-retrieval mechanism is broadly applicable across different model architectures and confirms the generalizability of our approach beyond specific model families.
@@
Appendix C Statistical Significance
We compute the 9595% confidence intervals for our main results in Table 1 and  2 using the bootstrap resampling method. On all datasets except the Bamboogle dataset, we observe that the performance of CoRAG-8B is significantly better than the baselines.
Appendix D Prompts